Find the linear velocity of the tip of the second hand of the clock if the hand is 30 mm long.
Found 2 solutions by TimothyLamb, Alan3354:
Answer by TimothyLamb(4379)

C = 2*pi*r = 2*pi*(30/1000 meters)~= 0.1884955 meters
---
length of 1-second segment on clock face:
C/60 ~= 3.1415927E-3 ~= 0.0031415 meters
---
s = d / t
---
linear speed of the clock's second hand:
(C/60)meters /(1 second) ~= 0.0031415 m/sec
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by Alan3354(69443)

Find the linear velocity of the tip of the second hand of the clock if the hand is 30 mm long.
------
In 1 revolution, it moves 2*pi*30 mm = 60pi mm
That takes 1 minute
Speed = 60pi mm/minute
=~ 188.5 mm/minute